OpenStudy (anonymous):

what is h + 8 = 11???

jimthompson5910 (jim_thompson5910):

You want to undo the +8, so subtract 8 from both sides to isolate h

OpenStudy (anonymous):

ok well how do i do that?

jimthompson5910 (jim_thompson5910):

just subtract 8 from 8 on the left side (that gives you 0 and it goes away) then subtract 8 from 11 on the right side (11 - 8 becomes ???)

OpenStudy (anonymous):

-3

jimthompson5910 (jim_thompson5910):

11 - 8 is not -3

jimthompson5910 (jim_thompson5910):

but you are close

OpenStudy (anonymous):

On any equation, if you do the same thing on both sides, it remains valid. In this case you want to eliminate the 8, so you can subtract 8 from both sides of the equation. you end up with h+8-8=11-8 on the left side the +8 and -8 eliminate each other, and on the right side you get a value, so the end result is: h=?

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

3

jimthompson5910 (jim_thompson5910):

good, so h = 3 is the answer

jimthompson5910 (jim_thompson5910):

Check: h + 8 = 11 3 + 8 = 11 ... Plug in h = 3 11 = 11 So the answer is verified

jimthompson5910 (jim_thompson5910):

h + 8 = 11 is basically asking "what number plus 8 equals 11?" The answer to that question is 3 since 3+8 = 11

OpenStudy (anonymous):

???????

jimthompson5910 (jim_thompson5910):

I'm just showing that 3 is indeed the correct answer

OpenStudy (anonymous):

o so i was right?

jimthompson5910 (jim_thompson5910):

your second answer, yes h = 3 is the answer

OpenStudy (anonymous):

ohh lol blonde moments haa!

OpenStudy (anonymous):

yes its 3

jimthompson5910 (jim_thompson5910):

lol that's perfectly fine

OpenStudy (anonymous):

i have a nother q

jimthompson5910 (jim_thompson5910):

ok what is it

OpenStudy (anonymous):

what is -6 + t = -14?

OpenStudy (anonymous):

message me for help

jimthompson5910 (jim_thompson5910):

-6 + SOME NUMBER = -14 is the same as saying -6 + t = -14 We just replace that "SOME NUMBER" with a variable t

OpenStudy (anonymous):

k

jimthompson5910 (jim_thompson5910):

to isolate t, we have to undo the -6, so we have to add 6 to both sides

jimthompson5910 (jim_thompson5910):

On the left side: -6+6 =0, so this goes away On the right side: -14 + 6 becomes ???

OpenStudy (anonymous):

so it would be 6+6?

OpenStudy (anonymous):

-

OpenStudy (anonymous):

noooooo

jimthompson5910 (jim_thompson5910):

close, -6+6

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

im sooo confuseddd

OpenStudy (anonymous):

message me

jimthompson5910 (jim_thompson5910):

on the left side, you add 6 to -6 and that will give you zero

jimthompson5910 (jim_thompson5910):

it will go away

jimthompson5910 (jim_thompson5910):

you do the same thing to the other side to get -14 + 6 = ???

OpenStudy (anonymous):

20

jimthompson5910 (jim_thompson5910):